OpenStudy (zenmo):

A small software company invests $25,000 to produce a software package that will sell for $69.95. Each unit can be produced for $45.25. How many units must be sold to make a profit of $100,000?

OpenStudy (shaik0124):

see profit for 1 unit will be profit=selling price-cost price =69.95-45.25 =24.70

OpenStudy (shaik0124):

for getting a profit of 24.70 u have to sell 1 unit now calculate for 100000

OpenStudy (zenmo):

$100,000=24.7x? something like that? This problem involves the method of substitution of linear and nonlinear system equations if that helps.

OpenStudy (shaik0124):

yeah it's correct

OpenStudy (shaik0124):

x=100000/24.7

OpenStudy (zenmo):

I'm getting 4048.58, but the solution says 5061 units

OpenStudy (zenmo):

for Part A, i had to find the break even point. Y=25,000 + 45.25x, Y=69.95x 25,000+42.25x=69.95x 25,000=24.7x x=1013 units.

OpenStudy (shaik0124):

is 69.95 the selling price for 1 unit??

OpenStudy (zenmo):

P=R-C R=69.95x, C=25,000+45.25x 100,000=69.95x-(25,000+45.25x) 100,000=69.95x-25,000-42.25x 125,000=24.7x x=5060.7287

OpenStudy (shaik0124):

for $25000 number of units produced =25000/45.25 =552.48618 units if we sell each unit at 69.95 then total profit for 552.48618 units is 552.48618*69.95=38646.408

OpenStudy (shaik0124):

what is this r=69.95x??

OpenStudy (shaik0124):

what does x mean in that

OpenStudy (zenmo):

I actually found the answer with the previous reply.

OpenStudy (shaik0124):

ok I know that.Just wanna know how did they solve

OpenStudy (zenmo):

oh no, I solved it. I posted it above. I only gave u a badge for sticking it out with me.

OpenStudy (shaik0124):

ok thanks